Find the y-intercept, the equation of the axis of symmetry, and the x-coordinate of the vertex.
Consider f(x)=x^2+2x-3

Answer by Alan3354(69443)

Find the y-intercept, the equation of the axis of symmetry, and the x-coordinate of the vertex.
Consider f(x)=x^2+2x-3
-------------------------
equation of the axis of symmetry is x = -b/2a
x = -2/2 = -1
-----
x coordinate = -1